An ideal Otto cycle has a compression ratio of 9. At the beginning of the compression process, air is at 100 kPa and 17°C, and 900 kJ/kg of heat is transferred to air during the constant-volume heat-addition process. Taking into account the variation of specific heats with temperature, determine (a) the pressure and temperature at the end of the heat-addition process, (b) the net work output, (c) the thermal efficiency, and (d) the mean effective pressure for the cycle .
Answer #1

P1           =             100 kPa

T1           =             17+273 =             290 K

h             =             900 kJ/kg

r              =             9

Stage 1

V1           =             (R*T1)/P1

                =             0(.287*290)/100

                =             0.83 m3/kg

Stage 2

Compression ratio

r              =             V1/V2

V2           =             0.83/9

                =             0.092 m3/kg

Vr1         =             676.1

U1          =             206.91 kJ/kg

Process 1-2

To find specific volume Vr

(V2/V1)                =             (Vr2/Vr1)

Vr2         =             Vr1/9

                =             676.1/9 =             75.12

So,

From gas tables

T2           =             680 k

U2          =             496.62 kJ/kg

P2           =             (R*T2)/V2

                =             (0.287*680)/0.092

                =             2121.3 K

Process 1 to 2

Q-W       =             m(U2-U1)

if m=1

W            =             U1-U2

                =             206.91-496.62

                =             -289.71 kJ/kg

Stage 3

V3           =             0.1133 m3/kg      =             V2

Process 2-3

h             =             750 kJ/kg             =             U3-U2

So, U3   =             900+496.62

                =             1396.62 kJ/kg

From gas tables

T3           =             1700 K

Vr3         =             4.761

So,

P3           =             (R*T3)/V3

                =             (0.287*1700)/0.092

                =             5303.2 kPa

Stage 4

V4           =             V1           = 0.83 m3/kg

(V4/V3)                =             (Vr4/Vr3)

Vr4         =             4.761*9

                =             42.849

From gas table

T4           =             840 K

U4          =             624.95 kJ/kg

P4           =             (R*T4)/V4

                =             (0.287*840)/0.83

                =             290.45 kPa

Process 3-4

Q-W       =             m(U4-U3)

For Q=0 and m=1

W            =             U3-U4

                =             1396.62-624.95

                =             771.67 kJ/kg

Process 4-1

W            =             U4-U1

                =             624.95 - 206.91

                =             418.04 kJ/kg

(b) So, net work

Wnet     =             Work in                -Work out

                =             900 - 418.04

                =             481.96 kJ/kg

(c) Efficiancy       =             481.96/900

                =             53.5%

(d) MEP                =             Wnet/(V1*(1-(1/r))

                =             481.96/(0.83*(1-1/9))

                =             653.25 kPa